OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 |x − 4| = 6 |3x + 12| = 18 Equation #1 and Equation #2 have the same number of solutions. Equation #1 has more solutions than equation #2. Equation #1 has fewer solutions than equation #2. None of the statements above describe the number of solutions to equations shown.

OpenStudy (cwrw238):

x - 4 = 6 or x - 4 = -6 second equation 3x + 12 = 18 or 3x + 12 = -18 so why do you think?

OpenStudy (cwrw238):

* lol what do you think?

OpenStudy (anonymous):

umm im not sure

OpenStudy (cwrw238):

they both have 2 solutions

OpenStudy (cwrw238):

the modulus signs | | indicate the values inside them make the values positive if they are negative or positive

OpenStudy (anonymous):

so a??

OpenStudy (cwrw238):

first option is correct

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thxs